OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 5x + 3y = 8 y = −2x + 1

OpenStudy (anonymous):

I would use the substitution method because you already have "y" by itself. 5x + 3y = 8 y = -2x + 1 Now substitute -2x + 1 in for y in the first equation 5x + 3(-2x + 1) = 8 5x - 6x + 3 = 8 -x = 8 - 3 -x = 5 x = - 5 Now substitute -5 in for x in the second equation y = -2x + 1 y = -2(-5) + 1 y = 10 + 1 y = 11 check.... 5x + 3y = 8 5(-5) + 3(11) = 8 -25 + 33 = 8 8 = 8 (correct) ANSWER : Solving through the substitution method : x = -5 and y = 11

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome